# No Language is an Island: Unifying Chinese and English in Financial Large Language Models, Instruction Data, and Benchmarks

### v2

@@ 3.2.1 Chinese Financial Datasets
We mainly focus on Chinese language proficiency. Thus, for Chinese, we provide two data types distinguished by their source of the labels: DLC and DLE, which encompass classification (ZH-CLS), extraction (ZH-EXT), prediction (ZH-PRE) and generation (ZH-GEN) task in financial NLP.
@@
For ZH-CLS, our dataset covers essential tasks such as sentiment analysis (FinSA), semantic matching (FinSM), news classification (FinNC), negative judgment (FinNJ), answer selection (FinAS) and relationship extraction (FinRE). 1) Sentiment Analysis. It is the process of analyzing and labeling the sentiment expressed in financial texts as "Negative(消极)","Neutral(中性)" and "Negative(消极)" (Sohangir et al., 2018). We used the FE (Lu et al., 2023) dataset sourced from Stockbar and Xueqiu websites. We also utilized the StockB 333https://huggingface.co/datasets/kuroneko5943/stock11 dataset, which focuses on the finance in among data from ten domains. 2) Semantic Matching. It determines if financial texts or sentences are semantically similar ("是(YES)" or "否(NO)"), using two datasets: the Bank Question Corpus (BQC) (Chen et al., 2018) from Chinese bank customer logs and the Ant Financial Question Matching Corpus (AFQMC) (Xu et al., 2020) from the Alipay competition, both utilizing binary classification for match identification. 3) Negative Judgment. It focuses on identifying finance and economics-related negativity, distinct from general financial sentiment analysis. Leveraging the FinNSP dataset (Lu et al., 2023) sourced from social media, it uses "是(YES)" to label data containing negative information and "否(NO)" for those without. 4) News Classification. It is to categorize news into 2 or multiple-class label, such as "股票市场(stock market)", "外汇(foreign exchange)", "宏观经济(macroeconomics)", and more. We use the FinNL (Lu et al., 2023) dataset, sourced from financial news websites like Sina Finance, Tencent Finance, Phoenix Finance, 36Kr, and Huxiu. It includes binary (NL) and multi-class (NL2) labeled data. 5) Answer Selection. FinEval (Zhang et al., 2023b) and and AntGroup evaluates financial reasoning by selecting the correct option ("A", "B", "C", "D) in a financial context. The dataset 444https://github.com/alipay/financial_evaluation_dataset includes 4,661 questions covering finance, economy, accounting, and certifications. We’ve compiled data for ten key finance subjects as FinEvalF. 6) Relationship Extraction. It aims to provide the types of connections, dependencies, and associations between entities. We gathered the FinRE dataset (Lu et al., 2023) and named it RE, which includes financial news articles along with head-tail entity pairs, covering 44 relationship categories (e.g., "合作(cooperation)", "持股(shareholding)", etc.), including the "未知(unknown)" relation.
@@
For ZH-EXT, our dataset includes question answering (FinQA), named entity recognition (FinER), and event detection (FinED). 1) Question Answering. It aims to extract accurate answers from the given text to answer questions related to an event mentioned in the text. We have collected a dataset named QA to differentiate from the English dataset FinQA (Lu et al., 2023) with the same name. 2) Named Entity Recognition(NER). It aims to identify entity names and types labeled as individuals ("个人(PER)"), organizations ("组织(ORG)"), or locations ("地点(LOC)") in financial text. We utilize the the comprehensive NER dataset, referred to as CNER in (Jia et al., 2020). 3) Event Detection. It involves recognizing and comprehending specific incidents beyond industry categorization. We collect four different financial challenge datasets 555https://www.biendata.xyz/competition/{ccks_2019_4,ccks_2020_4_1,ccks_2021_task6_2,ccks2022_eventext} from China Conference on Knowledge Graph and Semantic Computing (CCKS), named as 19CCKS, 20CCKS, 21CCKS and 22CCKS, respectively. Unlike similar work (Lei et al., 2023) that focuses on detecting 254 different event types in four categories (industry-related, macrography-related, market-related, and company-related), our collected and organized data further includes the task of identifying causal relationships between events (原因事件类型/Cause event type→结果事件类型/Result event type).
@@
For ZH-PRE, we focus on the Stock Prediction (FinSP) task. It involves categorizing stocks into three categories ("表现不佳(underperforming)", "跑赢大盘(outperforming the market)", "中性(neutral)") by combining stock data and financial news. We collected the StockA dataset (Zou et al., 2022), features stock-specific news and factors for China’s A-shares market, including minute-level price history, setting it apart from similar datasets (Xu and Cohen, 2018; Zhou et al., 2021) by offering comprehensive data on individual stocks.
For ZH-GEN, we include the Text Summarization (FinTS) task. The goal is to distill intricate financial details into succinct summaries. We curated the FinNA dataset (Lu et al., 2023), known as NA, by extracting information from research reports and specifically honing in on their abstracts.
3.2.2 English Financial Datasets
@@
For English, we utilize datasets from PIXIU Xie et al. (2023b), a comprehensive source for English financial data tailored for FinLLMs. The dataset includes five task in financial NLP: classification (EN-CLS), prediction (EN-PRE), extraction (EN-EXT), reasoning (EN-REA) and data-out-of-field (dof) task not used for training LLMs (EN-DOF).
@@
For EN-CLS, our dataset consists of the following three tasks: sentiment analysis (FinSA) , headline classification (FinHC) and credit classification (FinCC). 1) Sentiment Analysis. We used the Financial Phrase Bank (FPB) (Malo et al., 2014) and FiQA-SA (Maia et al., 2018) dataset. FPB comprises a human-annotated phrase bank for financial and economic texts, while FiQA-SA primarily originates from social data, including microblogs, news. 2) Headline Classification. We utilize the Headlines dataset (Sinha and Khandait, 2021), containing news headlines from 2000 to 2019 related to "gold" and annotated with categories like price movements (up, down, stable), historical and future prices, general trends, and asset comparisons. 3) Credit Classification. It involves assessing a customer’s credit status based on a set of attributes and assigning a credit label of "good" or "bad" to the customer. We employed the two datasets, German (Hofmann, 1994) and Australian (Quinlan, ).
@@
For EN-PRE, we approach Stock Prediction as a binary classification task (Soun et al., 2022), utilizing social media, financial news, and historical stock price data from the past five days to forecast the next-day close price change rate. Classification criteria are: 1) Close price change rate > 0.50% is labeled "Rise"; 2) Close price change rate < -0.50% is labeled "Fall"; 3) Other cases are labeled as "Hold." We utilize three common English datasets: ACL18 (Xu and Cohen, 2018), CIKM18 (Wu et al., 2018), and BigData22 (Soun et al., 2022).
@@
For EN-EXT, we include the NER task. We use the CoNLL-2003 dataset (Alvarado et al., 2015) named as NER including sentences from public financial agreements via U.S. SEC filings. It includes four annotated entity types: Location ("LOC"), Organisation ("ORG"), Person ("PER"), and Miscellaneous ("MISC").
@@
For EN-REA, We include the Question Answering task distinct from Chinese, additionally incorporating stock time series and tables to infer answers. We utilized the datasets FinQA (Chen et al., 2021) and ConvFinQA (Chen et al., 2022). FinQA comprises expert-written financial reports of S&P 500 companies, sourced from FinTabNet (Zheng et al., 2021). ConvFinQA expands on FinQA, allowing multi-turn conversations.
Additionally, we introduce EN-OFT including NER, text summarization (FinTS) and hawkish-dovish classification (FinDC) task. 1) Named Entity Recognition. In NER task, we utilize the FINER-ORD (Shah et al., 2023b) dataset, which provides labels for entities such as "PER", "LOC", and "ORG". 2) Text Summarization. For this task, we use two datasets: ECTSUM (Mukherjee et al., 2022) and EDTSUM (Zhou et al., 2021). ECTSUM involves extracting key sentences from texts, while EDTSUM focuses on generating fitting news headlines through abstractive summarization. 3) Hawkish-dovish Classification. FinDC aims to classify sentences from monetary policy texts into a "hawkish" or "dovish" stance, unlike FinSA, using the FOMC (Shah et al., 2023a) dataset.
3.2.3 Translation Datasets
@@
To enhance the model’s bilingual and cross-lingual capabilities, we utilized ChatGPT to translate 8 English datasets (ZH-TRA) spanning 4 tasks: FinSA, FinSP, FinHC, and FinQA, into Chinese. We fully translated datasets for text-centric tasks like FinSA and FinHC, while translating only the validation sets for the more intricate FinSP and FinQA tasks. The English FinSA datasets FPB and FiQA-SA were rendered as CFPB and CFiQA-SA in Chinese, featuring sentiment labels: "消极(Negative)", "中性(Neutral)", and "积极(Positive)". The Headlines dataset was transformed into CHeadlines in Chinese, with labels "是(Yes)" and "否(No)". For QA tasks like EnQA and ConvFinQA, we tailored prompts to preserve vital information such as tables, stock names, resulting in new Chinese datasets named CEnQA and CConvFinQA. In FinSP tasks like ACL18, CIKM18, and BigData22, we employed regular expressions to extract text containing task prompts and stock descriptions. Subsequently, by using human-crafted translation prompts, we translated the text while retaining stock details like stock tickers and historical stock prices, yielding translated Chinese datasets named CACL18, CCIKM18, and CBigData22, respectively. To maintain data quality, we removed error-prone or untranslatable data. Our translation method example is available in Appendix Figure 5.
4 ICE-FIND: Bilingual Financial Instruction Data
Base on the raw datasets, we further construct our bIlingual Chinese-English Financial InstructioN Dataset (ICE-FIND), covering 13 tasks and 36 datasets with 604k data samples in total, as detailed in Appendix Table 7. This dataset compilation combined raw data with expertly devised prompts, generating 30-40 unique prompts for each dataset. However, it is important to note that expert-authored prompts may not always accurately reflect the quality of the generated results. We assessed these prompts through human evaluations 666Examples and thier translations are shown in Appendix Figure 8 focusing on accuracy, naturalness, and informativeness using a 1-3 scale, to identify the most effective ones. These prompts were tested on various LLMs such as Bing Chat, ChatGLM using the ChatALL platform Sun (2024), showing satisfactory performance in most cases.
@@ 4 ICE-FIND: Bilingual Financial Instruction Data
To handle prompt variations across LLMs, we adopte two strategies. For English tasks, we match each data with all prompts except for FinQA and ConvFinQA. For Chinese tasks, we randomly select a prompt for each data due to the larger scale of the raw data. Finally, we convert the raw datasets into instruction-tuning samples by by restructuring expert prompts 777Examples are presented in Appendix Table 10, input texts, and responses. The instruction data was structured in JSON format (Figure 3). {data_id} represents the ID of each data, {prompt} is the task-specific prompt, {conext} refers to the input data, {Your answer} denotes the expected answer, and {Your label} represents the truly answer.
5 ICE-INTERN: Bilingual Financial Large Language Model

### v3

@@ 3.1.1. Raw Data Integration
(A) Chinese Financial Datasets. Our primary focus is on Chinese language proficiency, with the cleaned and processed datasets can be categorized into two data types: DLC and DLE, which encompass classification (ZH-CLS), extraction (ZH-EXT), prediction (ZH-PRE) and generation (ZH-GEN) financial NLP tasks.
@@
For ZH-CLS, it covers essential tasks like sentiment analysis (FinSA), semantic matching (FinSM), news classification (FinNC), negative judgment (FinNJ), answer selection (FinAS) and relationship extraction (FinRE). 1) Sentiment Analysis. It is the process of analyzing and labeling the sentiment expressed in financial texts as "Negative(消极)","Neutral(中性)" and "Negative(消极)" (Sohangir et al., 2018). This task integrates the FE (Lu et al., 2023) dataset from Stockbar and Xueqiu websites, along with the StockB (Wang, 2023) dataset that covers finance across ten domains. 2) Semantic Matching. It determines whether financial texts or sentences are semantically similar ("是(YES)" or "否(NO)"), using two datasets: the Bank Question Corpus (BQC) (Chen et al., 2018) from Chinese bank customer logs and the Ant Financial Question Matching Corpus (AFQMC) (Xu et al., 2020) from the Alipay competition. 3) Negative Judgment. It focuses on identifying finance and economics-related negativity, as opposed to general financial sentiment analysis. Leveraging the FinNSP dataset (Lu et al., 2023) sourced from social media, it uses "是(YES)" to label data with negative information and "否(NO)" for those without. 4) News Classification. It involves classifying news into two or more class label, such as "股票市场(stock market)", "外汇(foreign exchange)", "宏观经济(macroeconomics)", and more. This task employs the FinNL (Lu et al., 2023) dataset, sourced from financial news websites, containing both binary (NL) and multi-class (NL2) labeled data. 5) Answer Selection. It entails selecting the correct option ("A", "B", "C", "D") within the financial context to evaluate financial reasoning and decision-making skills. AntGroup’s FinEval (Zhang et al., 2023a; AntNathangq, 2023) includes 4,661 questions covering finance, economy, accounting, and certification. We compile the available financial subject data into FinEvalF. 6) Relationship Extraction. It aims to provide the types of connections, dependencies, and associations between entities. We process the FinRE dataset (Lu et al., 2023) into RE, comprising financial news articles with head-tail entity pairs, spanning 44 relationship categories (e.g., "合作(cooperation)", "持股(shareholding)", etc.).
@@
For ZH-EXT, it includes the question answering (FinQA), named entity recognition (FinER), and event detection (FinED) tasks. 1) Question Answering. It aims to extract precise answers from financial text for questions about pertinent events. We gather a dataset named QA (Lu et al., 2023) to differentiate its name from the English FinQA dataset. 2) Named Entity Recognition(NER). It aims to identify entity names and types labeled as individuals ("个人(PER)"), organizations ("组织(ORG)"), or locations ("地点(LOC)") in financial text. We utilize the the comprehensive NER dataset (Jia et al., 2020) called CNER. 3) Event Detection. It involves recognizing and comprehending specific incidents beyond industry categorization. This task integrates the 19CCKS, 20CCKS, 21CCKS, and 22CCKS datasets (CCKS, 2022) from the China Conference on Knowledge Graph and Semantic Computing (CCKS). Unlike CFGPT (Li et al., 2023a) that focuses on detecting 254 different event types across four categories, our task additionally involves identifying causal relationships between events ("原因事件类型(Cause event type)"→"结果事件类型(Result event type)").
@@
For ZH-PRE, it centers on the Stock Prediction (FinSP) task, which entails categorizing stocks into three categories ("表现不佳(under-performing)", "跑赢大盘(outperforming the market)", "中性(neutral)") by combining stock data and financial news. This task utilizes the StockA dataset (Zou et al., 2022), showcasing stock-specific news and factors within China’s A-shares market. It stands out from similar datasets (Xu and Cohen, 2018; Zhou et al., 2021) by delivering thorough data on individual stocks.
For ZH-GEN, it encompasses the Text Summarization (FinTS) task, aiming to condense complex financial information into concise summaries. The FinNA dataset (Lu et al., 2023), referred to as NA for this task, consists of financial news articles along with their abstracts.
@@
(B) English Financial Datasets. For English, we integrate datasets from PIXIU (Xie et al., 2023) tailored for English-centric FinLLMs. It includes five tasks in financial NLP: classification (EN-CLS), prediction (EN-PRE), extraction (EN-EXT), reasoning (EN-REA).
@@
For EN-CLS, it consists of three tasks: sentiment analysis (FinSA), headline classification (FinHC) and credit classification (FinCC). 1) Sentiment Analysis. It uses the Financial Phrase Bank (FPB) (Malo et al., 2014) and FiQA-SA (Maia et al., 2018) datasets. FPB comprises a human-annotated phrase bank for financial and economic texts, while FiQA-SA mainly comes from social data. 2) Headline Classification. It utilizes the Headlines (Sinha and Khandait, 2021) dataset, comprising gold-related news headlines from 2000 to 2019 annotated with nine categories like "price up”, "price down", "price stable”, "future price", and "asset comparisons". 3) Credit Classification. It aims to assess a customer’s credit status using a set of attributes and assigning a credit label of "good" or "bad". This task employs the German (Hofmann, 1994) and Australian (Quinlan, [n. d.]) datasets.
@@
For EN-PRE, it treats Stock Prediction (FinSP) as a ternary classification task (Soun et al., 2022), utilizing social media, financial news, and historical stock price data to forecast the close price change rate. The classification criteria are as follows: 1) Close price change rate > 0.50% is labeled "Rise"; 2) Close price change rate < -0.50% is labeled "Fall"; 3) Other cases are labeled as "Hold". Three common datasets used are ACL18 (Xu and Cohen, 2018), CIKM18 (Wu et al., 2018), and BigData22 (Soun et al., 2022).
@@
For EN-EXT, it centers on the NER (FinER) task. We use the CoNLL-2003 (Alvarado et al., 2015) dataset, named NER, for annotating financial agreements from U.S. SEC filings into location ("LOC"), organization ("ORG"), person ("PER"), and miscellaneous ("MISC") entities.
@@
For EN-REA, it differs from the Chinese Question Answering (FinQA) task in that it integrates stock time series and tables to infer answers. Two datasets used are FinQA (Chen et al., 2021) and ConvFinQA (Chen et al., 2022). FinQA consists of expert-annotated Q&A pairs derived from financial reports of S&P 500 companies within the FinTabNet dataset (Zheng et al., 2021). ConvFinQA extends FinQA to allow for multi-turn conversations.
(C) Chinese Translation Datasets. To enhance the model’s cross-lingual capabilities and explore the performance influence of incorporating translated data, we create new English-to-Chinese Translation (ZH-TRA) Financial NLP datasets, covering 4 specific financial tasks: sentiment analysis (FinSA), headlines classification (FinHC), question answering (FinQA) and Stock Prediction (FinSP).
@@
Specifically, we utilize ChatGPT 3.5-Turbo-1106 to translate the 8 English datasets involved in these specific tasks into Chinese. Notably, we perform complete translations (including training, validation, and testing sets) for the only-text modality datasets within the FinSA and FinHC tasks. Conversely, for FinSP and FinQA, which incorporate additional sequence and tabular modalities, we focus on translating the only testing sets. In FinSA, the English datasets FPB and FiQA-SA are translated into Chinese as CFPB and CFiQA-SA, while in FinHC, the English Headlines dataset is translated into Chinese as CHeadlines. In FinQA, we specifically tailor translation prompts for the EnQA and ConvFinQA English datasets to preserve key information like tables and stock names, creating the Chinese CEnQA and CConvFinQA datasets. In FinSP, we utilize regular expressions to extract text containing task prompts and stock descriptions. With manually crafted translation prompts, we translate the text descriptions while retaining stock details like stock tickers and historical stock prices, resulting in the Chinese datasets named CACL18, CCIKM18, and CBigData22, respectively.
To balance data integrity and machine translation quality, we eliminate the flaw data as follows: 1) Expert review of translation discrepancies in language structures and specialized terms (Participant details as outlined in Section 3.1.2) . 2) ChatGPT-turbo’s constraint of 4096 tokens, unable to process excessively long texts with time-series data. An example of the translation method used for the Bigdata22 dataset in FinSP is depicted in Appendix Figure 3.
(D) Out-of-Training Financial Datasets. A factor overlooked by nearly all existing methods is the introduction of out-of-training datasets (not used for model fine-tuning) to explore their robustness and generalization ability. Thus, we introduce the EN-OFT financial NLP task including NER (FinER), text summarization (FinTS) and hawkish-dovish classification (FinDC), which are not used during the training process. 1) Named Entity Recognition. It utilizes the FINER-ORD (Shah et al., 2023b) dataset, which provides three labels for entities such as "PER", "LOC", and "ORG". 2) Text Summarization. Two datasets ECTSUM (Mukherjee et al., 2022) and EDTSUM (Zhou et al., 2021), are utilized for this task. ECTSUM involves extracting key sentences from texts, while EDTSUM focuses on generating fitting news headlines through abstractive summarization. 3) Hawkish-dovish Classification. Unlike FinSA, it involves classifying sentences from monetary policy texts into a "hawkish" or "dovish" stance using the FOMC (Shah et al., 2023a) dataset.
3.1.2. Instruction Construction
Based on raw data, we further build bIlingual Chinese-English Financial InstructioN Dataset (ICE-FIND), covering 20 specific tasks and 40 datasets with a total of 604k data samples, as detailed in Table 2. Building ICE-FIND requires integrating diverse prompts into raw data. We convene domain experts to write 30-40 unique prompts for each dataset. Examples of prompts and their translations are shown in Appendix Figure 5. However, it is crucial to acknowledge that expert-authored prompts may not consistently yield high-quality results. Thus, we undertake a manual evaluation of the prompts to ensure their effectiveness.
Specifically, the evaluation panel consists of 8 graduate students and 2 undergraduate students, all proficient in both Chinese and English, with backgrounds in finance domain. Each participant is provided with a questionnaire comprising 3-4 prompts (across 10 rounds) for quality assessment, covering three metrics: accuracy, naturalness, and informativeness. The questionnaire options include: "Strongly Disagree(0)", "Disagree(1)", "Agree(2)", "Strongly Agree(3)". Participants test the prompts with a randomly selected LLM on the ChatALL (Sun, 2023) platform to rate each prompt from 0 to 3 across all metrics. ChatALL contains various LLMs, such as Bing Chat (Mcirosoft, 2022) and ChatGLM (Du et al., 2021). We retain the prompts with an average score above 2, ensuring adaptability to most cases. Examples of human evaluation are presented in Appendix Table 9-10.
@@ 3.1.2. Instruction Construction
To address prompt variation in LLMs, we adopt two strategies. For English tasks, we pair each data with all prompts except for FinQA and ConvFinQA. For Chinese tasks, given the larger raw data, one prompt is randomly selected for matching each data. Finally, we convert the raw datasets into instruction-tuning samples by restructuring expert prompts, input texts, and responses. Instruction tuning samples adheres to the structured template as follows.
| {Id:[data_id], |
| {Instruction:[prompt]Input:[text]{Output:[response]} |
which integrates human-designed instructions with input texts and their corresponding responses. [data_id] represents the ID of each data, [prompt] is the task-specific prompt, [text] refers to the input text, [response] denotes the expected answer.
3.2. ICE-INTERN: Bilingual Financial Large Language Model
Upon evaluating bilingual performance of various backbone LLMs on ICE-FLARE, we explore using InternLM-7B as backbone model with ICE-FIND fine-tuning to develop a bilingual FinLLM named ICE-INTERN-7B. To assess the impact of various data types on fine-tuning ICE-INTERN, we gradually incorporate the DLC, DLE, DTT and DTE datasets during the fine-tuning process, creating ICE-INTERN variants named ICEdlc-7B (ZH), ICEdle-7B (ZH), ICEdtt-7B (ZH), and ICEfull-7B (ZH & EN), respectively. To optimize memory usage and reduce training times, we introduce an advanced fine-tuning technique called QLoRA (Hu et al., 2021). The same fine-tuning hyper-parameters set for all models are described follows. Specifically, the training sequences are uniformly cut to a length of 2048 tokens. Then, we use the AdamW (Kingma and Ba, 2014) optimizer with an initial learning rate of 5e-5 on a weight decay 1e-5 for all parameters. Additionally, the warm-up steps are set to 1% of the total training steps. Finally, we fine-tune the model with 1 epoch in batch size of 24 on 8 NVIDIA HGX A100 SXM4 GPUs. It is worth noting that our ICEdtt-7B and ICEfull-7B models can be used for language-specific ablation studies, while combining the ICEdle-7B and ICEdtt-7B models enables a deeper exploration of how task-specific Chinese translated datasets can enhance the model’s bilingual capabilities.
3.3. ICE-FLARE: Bilingual Financial Evaluation Benchmark
Based on ICE-FIND, we propose the ICE-FLARE bilingual evaluation benchmark. ICE-FLARE can evaluate the bilingual capabilities of LLMs in the financial domain. The comparison of existing benchmarks ( [a] CFBenchmark (Lei et al., 2023), [b] FinCUGE (Chen et al., 2023), [c] Fineval (Zhang et al., 2023a), [d] CGCE (Zhang et al., 2023b), [e] CFinBench (Nie et al., 2024b), [f] FLUE (Shah et al., 2022),[g] FLARE (Xie et al., 2023), [h]FinanceBench (Islam et al., 2023), [i]BizBench (Koncel-Kedziorski et al., 2023), [j]FinBen (Xie et al., 2024)) with ICE-FLARE is detailed in Table 3. Unfortunately, they are all designed exclusively for either Chinese or English. From Table 3, it is evident that ICE-FLARE, which introduces bilingual capabilities, significantly surpasses these benchmarks on task diversity and data scalability. Despite FinCUGE having a higher number of financial evaluation tasks, the difference in dataset size compared to ICE-FLARE is nearly 8 times. These public benchmarks have provided valuable references and inspiration for constructing the test set, making ICE-FLARE more comprehensive in terms of task coverage, metric diversity, and sample quantity. However, they still face challenges in evaluating the bilingual capabilities in Chinese-English and lack DOT tasks (e.g., FinDC, FinST, FinER) to test the generalization ability of the FinLLMs. In conclusion, ICE-FLARE provides a bilingual evaluation benchmark that covers a sufficient range of diversity tasks and more comprehensive measurement metrics, which is crucial for improving the performance and linguistic capabilities of FinLLMs. An overview of ICE-FLARE, as shown in Figure 1.